OpenStudy (anonymous):

The profit of a company is modeled by the equation p(x) = 60x − 2x2 − 2, where p the profit is in thousands of dollars and x is the number of units sold in hundreds. How many units must the company sell to make the maximum profit? 1,000 1,500 2,000 2,500

OpenStudy (anonymous):

p(x)=60x-2x^2-2

OpenStudy (owlcoffee):

Well, in order to find the maximum profit, we would have to find, the functions maximum value. And to do that, we have to derivate the function: \[p(x)=60x-2x^2-2\] and derivating it: \[p'(x)=60-4x\] Then, a maximum is determined when the derivative is zero: \[60-4x=0\] \[x=15\]

OpenStudy (anonymous):

so what do I do with the 15?

OpenStudy (owlcoffee):

evaluate it in the original function P(x)

OpenStudy (anonymous):

plug in the 15s for x?

OpenStudy (owlcoffee):

yes

OpenStudy (anonymous):

60(15)-2(15)^2-2

OpenStudy (anonymous):

I didnt get nowhere near the answers

OpenStudy (owlcoffee):

what did you get?

OpenStudy (anonymous):

448

OpenStudy (owlcoffee):

yes, it's correct, but I gave you the answer, re-read the question.

OpenStudy (anonymous):

so its 1500

OpenStudy (owlcoffee):

correct